# spacing=nonfrench

### v1

@@ Abstract
Since the C++ bindings were deleted in 2008, the Message Passing Interface (MPI) community has revived efforts in building high-level modern C++ interfaces. Such interfaces are either built to serve specific scientific application needs (with limited coverage to the underlying MPI functionalities), or as an exercise in general-purpose programming model building, with the hope that bespoke interfaces can be broadly adopted to construct a variety of distributed-memory scientific applications. However, with the advent of modern C++-based heterogeneous programming models, GPUs and widespread Machine Learning (ML) usage in contemporary scientific computing, the role of prospective community-standardized high-level C++ interfaces to MPI is evolving. The success of such an interface clearly will depend on providing robust abstractions and features adhering to the generic programming principles that underpin the C++ programming language, without compromising on either performance and portability, the core principles upon which MPI was founded. However, there is a tension between idiomatic C++ handling of types and lifetimes, and, MPI’s loose interpretation of object lifetimes/ownership and insistence on maintaining global states.
Instead of proposing “yet another” high-level C++ interface to MPI, overlooking or providing partial solutions to work around the key issues concerning the dissonance between MPI semantics and idiomatic C++, this paper focuses on the three fundamental aspects of a high-level interface: type system, object lifetimes and communication buffers, also identifying inconsistencies in the MPI specification. Presumptive solutions can be unrefined, and we hope the broader MPI and C++ communities will engage with us in productive exchange of ideas and concerns.
Keywords:
Message Passing Interface C++ Concept-based Interface.
@@ 1 Introduction
Build on existing ideas. We consider, re-evaluate, formalize and extend upon ideas presented in existing language bindings (Section 2).
Derive a C++ representation of MPI’s object model. We discuss how to bridge the gap between MPI objects and C++, particularly how to handle their life cycle and formalize mutability in terms of constness (Section 3.1).
Enabling type safety in MPI. The MPI C API currently does not utilize type information provided by the language. In Section 3.2, we categorize C++ types and how these different classes can support type safe compatibility with MPI.
Contiguous ranges as first class communication buffers. Section 3.3 describes rules defined through C++-20 concepts, which specify which kind of C++ containers are directly supported by MPI for communication operations. This enables direct support for many STL containers and provides a powerful interface to derive more complex abstractions. We also describe how to handle data ownership which is necessary to conform to the best practices of resource management in modern C++ (namely, RAII) in Section 3.4.
Idiomatic error handling. We describe how C++ can enhance MPI error handling via compile-time checking and exceptions in Section 3.5.
These are the major aspects and principles to which a future C++ MPI interface should adhere. Rather than proposing a complete standardized interface, which risks repeating the shortcomings of the removed bindings, our primary goal is to define semantic guidelines and core conceptual interfaces that align idiomatic C++ with MPI. These concepts could serve as the foundation for a MPI standard side document. Standardizing concepts rather than concrete APIs offers the advantage of long-term maintainability and adaptability as both MPI and the C++ language evolve. However, a C++ interface could introduce additional quality-of-life features enhancing productivity and providing safer abstraction, but these often introduce overheads, conflicting with MPI’s performance-portability while increasing the complexity of a specification. Still, we outline major ideas and provide suggestions on how they could be aligned to our previously introduced design considerations. Particularly, we focus on serialization of more complex data types (Section 4.1), first steps towards automatic life cycle handling of more complex MPI data types with direct mappings to C++ (Section 4.2) and how to increase MPI programmer’s productivity by a sane set of defaults (Section 4.3).
In Section 5, we conclude our discussion with suggestions on how the MPI standard could support language interface designers (not limited to C++) by carefully extending the MPI specification.
@@ 2 Existing C++ interfaces over MPI
Since the removal of the “official” C++ MPI interface, a large number of third-party library interfaces has emerged, together with efforts to bridge the gap between MPI implementations, C++ and high-level performance portability frameworks such as Kokkos. We will briefly outline notable works with their distinctive features in the following, as they serve as inspirations and proofs of concept for the technical discussion in the rest of this paper.
@@
RWTH-MPI [9] is a C++ interface for MPI that supports contiguous STL containers as send/receive buffers and offers overloaded MPI operations with automatic parameter inference. For custom types, it can auto-generate MPI data types using the PFR library [19]. Dynamic-size types are unsupported. While it covers MPI 4.0, its bindings largely mirror the C interface with limited added abstraction or safety.
@@
Kokkos Comm [2] utilizes modern C++ language features in pursuit of more intuitive support of inter-node communication of Kokkos Views. Templates provide the needed flexibility to adapt to diverse data layouts and memory spaces.
@@ 3 High-level design considerations
We start by defining how MPI objects map to C++ (Section 3.1). Then we focus on how to model the actual data involved in communication, in terms of types (Section 3.2) and the actual memory involved (Section 3.3). We then discuss how combining C++’s ownership model and MPI to obtain additional memory safety (Section 3.4) and how to handle errors in an idiomatic way (Section 3.5).
@@ 3.1 Mapping the MPI object model to C++
For more than a decade, C++ MPI library designers have advocated for representing MPI objects as first-class C++ objects that act as proxies for the implementation-defined objects behind the handles [22]. Modern C++ idioms—particularly RAII (Resource Acquisition Is Initialization) [24, 16.5][25, E.6] and move semantics—naturally support resource management: MPI objects can be constructed via MPI_*_create in class constructors and released via MPI_*_free in destructors, with ownership safely transferable through move operations. However, the original MPI design deviates from this model because of its reliance on global state and implicitly managed global objects—an approach discouraged by the C++ Core Guidelines [25, I.3,I.22]. For example, global communicators such as MPI_COMM_WORLD are automatically initialized via MPI_Init and cannot be explicitly created or freed. Further, MPI prohibits reinitialization after MPI_Finalize, limiting modular or library-based usage patterns.
@@ 3.1.1 Sessions, Groups, and Communicators
MPI 4.0 addressed some of these limitations by introducing the Session Model, an alternative to the traditional World Model for process management. Although originally designed to support better isolation in multi-threaded and multi-component environments, the session model aligns well with modern C++ object-oriented design. In this model, no global communicators are predefined; instead, communicators must be explicitly created from process sets via user code. This makes ownership explicit and fully under user control, enabling clean integration with RAII-based designs as shown in Fig. 1. For example, communicators and groups are constructed and destroyed as part of their enclosing scope.
@@
The concept of constness is important for both the compiler and, more crucially, libraries, as it indicates variable mutability and restricts modifications in certain contexts. This applies straightforwardly to the data being communicated: For instance, a receiving buffer cannot be const, which already enhances the program’s safety, which we discuss in detail in Section 3.3. In a less obvious way, constness can also be applied to library objects of a MPI C++ interface, such as communicators, data types, and requests and operations on them, making it even more important to discuss here.
@@
We want to illustrate this for MPI communicators. Most operations leave a communicator in the same state as before the operation, which would naively allow them to be specified as const. Conversely, one could argue that posting a message is not the same as not having posted it; in the former case, the message can be received on the other side, while in the latter, a receive operation may hang. We think this is a strong argument to mark communication operations as non-const. But this can still be challenged for operations that send and receive in a single call (such as send_recv or collectives).  [^1] Note that this is a fundamental problem, not exclusive to const-qualified member functions; if the operations are free functions the questions will still stand regarding the constness of the communicator argument parameter.
To resolve these cases, we invoke a modern interpretation (post C++11): Since the communicator is likely undergoing internal mutation during any non-trivial operation, it is reasonable to conclude that most communication operations should be marked as non-const at the communicator level.
@@
auto size() const; /*most likely can be marked const in a reasonable implementation*/
@@
If we adhere to the conclusion that communication operations should not be marked const, an important and somewhat unexpected result follows: a communicator class should not have a duplicate mechanism that is const on the original communicator. Incidentally, there will be no canonical copy constructor (taking original communicator as a constant reference). This is consistent, since even communicators duplicated from each other cannot receive or complete operations initiated in another communicator. So two communicators are never copies of each other; they are at most alternative virtual fabrics for subsequent communication. Fig. 2 summarizes our discussions about the communicator class which also align with the interpretation of the B.MPI3 [7] interface.
@@ 3.2 Modeling and mapping types
Using template-metaprogramming, a C++ MPI interface can therefore deduce an MPI type directly from a data buffer (as defined in Section 3.3), in case its underlying value_type is fundamental. This approach is implemented by all major MPI C++ bindings [3, 7, 13, 26]. More complex types require explicit creation and a subsequent commit step, and have to be freed before MPI is finalized. To enable proper cleanup we again use RAII and represent data types as mpi::datatype objects, which support move construction and assignment and free the data type when the destructor is called.
To prevent users from using uncommitted data types in communication, we propose to encode the commit information as part of the type, similar to the approach of rsmpi [23]. The function mpi::commit takes an rvalue mpi::datatype object, and converts it to a mpi::committed_datatype as shown in Fig. 3. Accordingly, communication calls only accept data type objects of this type.
@@
double b;
char c;
std::array<int, 3> d;
@@
These types can be safely supported by automatically constructing the corresponding MPI datatype using a compile-time reflection mechanism. Although C++ currently lacks native language support for such reflection, concrete proposals are in place to introduce compile-time type reflection in C++26 [21]. In the meantime, third-party libraries such as Boost.PFR [19] can be used to implement this feature, although they have some limitations, for example for types using inheritance or private members.
Beyond fundamental and trivially copyable types, there remains a broad class of other (potentially user-defined) C++ types that do not have a direct or automatically derivable mapping to an MPI data type. For these types, the correspondence between memory layout and type semantics cannot be safely inferred by the prospective C++ MPI interface. Consequently, it becomes the user’s responsibility to explicitly define how the type is laid out in memory and how that layout corresponds to an appropriate MPI datatype. While the handling of fundamental and trivially copyable types can be automated to reduce programming errors and ease MPI development, a prerequisite for such automation is a clear conceptual model of the data involved in communication. The core abstraction in this model is the data buffer, which we introduce in the following Section 3.3.
@@ 3.3 Modeling Memory involved in MPI Communication
Particularly when using standard library containers such as std::vector<T> with the MPI C API, users have to access underlying raw pointers and sizes of the containers and pass them to MPI individually, while the actual vector object already comprises the whole send data context. This is both non-ergonomic to use and conflicts with the C++ Core Guidelines [25, I.13]. Most existing MPI C++ libraries therefore introduce support for a subset of standard library containers [3, 7, 9, 13, 26] as their core abstraction feature. However, this has certain shortcomings. First, this is often defined ad-hoc only including a fixed set of containers [9, 13]. Second, design decisions within the approaches of supporting (standard library) containers introduce hidden, additional overheads through memory allocations or additional MPI calls [3, 13, 26]. For example, in an MPI_Recv call, Boost.MPI resizes the container into which it receives the data to match the size of the incoming message. If the user does not want this, they must instead pass raw pointers again. The KaMPIng library internally invokes MPI_Probe when the user issues a receive operation without a count argument. Such unexpected overheads are not desired from the perspective of a standardized C++ interface.
@@
For each send or receive buffer involved in a (regular) communication call, MPI’s C API expects a pointer to a contiguous memory location, a count argument and an MPI data type. C++ already provides type information, and automatically matching types is crucial for type safety. A high-level standardizable C++ interface, can use type introspection to provide such type safety, which we detail in Section 3.2. For now, let us assume that we can always deduce an MPI data type from a provided container. We call this basic building block of MPI communication (memory location, count and type) a data buffer. In the following, we describe a minimal interface of data buffer objects in terms of C++ concepts.
@@
  • std::ranges::contiguous_range, that is its underlying memory has to be contiguous in memory. This captures MPI’s requirement that data which is communicated must reside in sequential storage  [^2] An exception to this requirement is the usage of MPI_BOTTOM which relies on absolute memory addresses to describe data and should be treated separately..
@@
An example of how these constraints can be expressed in plain C++ is given in Fig. 4, which defines the concept mpi::DataBuffer in terms of the three concepts described above. A convenient property of this data buffer definition is that many C++ standard library containers like std::vector or std::array storing fundamental C++ types such as char, int, double, ... already satisfy the mpi::DataBuffer concept. This also holds for lightweight, non-owning views such as std::span. Hence, they can be directly used in an MPI C++ interface adhering to these principles.
For more complex cases, for example, to communicate a container of a trivially_copyable custom C++ datatype, we first have to construct and commit such a datatype (see Section 3.2) and then pass it to the data buffer. For communicating data using complex derived datatypes without a one-to-one mapping to the underlying C++ type, we not only have to construct and commit an MPI datatype but also pass the number of corresponding elements to the data buffer. For these more complex cases, there is no direct C++ standard library support but an MPI C++ interface could easily provide a lightweight adapter with multiple constructors not only accepting a contiguous std::range, but also a count and a mpi::committed_datatype parameter. This also makes this interface extensible to support more complex containers from third-party libraries like Kokkos::View or thrust::device_vector. Figure 5 illustrates exemplary communication calls showcasing the different cases discussed above. Note that we do not suggest a concrete API or function signatures but rather want to show the underlying concept of using data buffers as an abstract description for communication data.
Orthogonal to the mpi::DataBuffer concept presented so far, we additionally have to model that data buffers describing send data have to be readable and data buffers for receive data have to be writable. This can be expressed using the (built-in) concepts, std::ranges::in/output_range (see Fig. 4).
@@
std::vector<MyType> v = ... // static custom type defined in Fig. 3
@@
Figure 5: Exemplary communication calls using the data buffer concept
Extending the data buffer concept
In irregular collective communication calls, MPI requires not just a single count parameter but a separate count for each MPI process (aka rank). Additionally, displacements can be specified for each rank.To reflect these semantic changes, we have to extend the mpi::DataBuffer concept as previously defined (Fig. 4) and define irregular mpi::DataBuffers. These buffers are still based on the concept std::ranges::contiguous and have to expose MPI datatype information. Additionally, they also have to expose size_v() and displacements() functions, which return the respective counts and displacement data. For brevity, we refrain from providing an example definition of this additional concept. Again, simple lightweight adapters can be provided for standard library and similar custom containers. Similarly, scalar arguments (such as a single int or bool in a reduction) can be easily supported via lightweight adapters, which implicitly return size $11$ and implement the data buffer concept (Fig. 5).

### v2

@@ Abstract
Since the C++ bindings were deleted in 2008, the Message Passing Interface (MPI) community has recently revived efforts in building high-level modern C++ interfaces. Such interfaces are either built to serve specific scientific application needs (with limited coverage to the underlying MPI functionality), or as an exercise in general-purpose programming model building, with the hope that bespoke interfaces can be broadly adopted to construct a variety of distributed-memory scientific applications. However, with the advent of modern C++-based heterogeneous programming models, GPUs and widespread Machine Learning (ML) usage in contemporary scientific computing, the role of prospective community-standardized high-level C++ interfaces to MPI is evolving. The success of such an interface clearly will depend on providing robust abstractions and features adhering to the generic programming principles that underpin the C++ programming language, without compromising on either performance or portability, the core principles upon which MPI was founded. However, there is a tension between idiomatic C++ handling of types and lifetimes and MPI’s loose interpretation of object lifetimes/ownership and insistence on maintaining global states.
Instead of proposing “yet another” high-level C++ interface to MPI, overlooking or providing partial solutions to work around the key issues concerning the dissonance between MPI semantics and idiomatic C++, this paper focuses on the three fundamental aspects of a high-level interface: type system, object lifetimes, and communication buffers, while also identifying inconsistencies in the MPI specification. Presumptive solutions can be unrefined, and we hope the broader MPI and C++ communities will engage with us in productive exchange of ideas and concerns.
@@ 1 Introduction
Build on existing ideas. We consider, re-evaluate, formalize and extend upon ideas presented in existing language bindings (Section˜2).
Derive a C++ representation of MPI’s object model. We discuss how to bridge the gap between MPI objects and C++, particularly how to handle their life cycle and formalize mutability in terms of constness (Section˜3.1).
Enabling type safety in MPI. The MPI C API currently does not utilize type information provided by the language. In Section˜3.2, we categorize C++ types and how these different classes can support type safe compatibility with MPI.
Contiguous ranges as first class communication buffers. Section˜3.3 describes rules defined through C++-20 concepts, which specify which kind of C++ containers are directly supported by MPI for communication operations. This enables direct support for many STL containers and provides a powerful interface to derive more complex abstractions. We also describe how to handle data ownership, which is necessary to conform to the best practices of resource management in modern C++ (namely, RAII) in Section˜3.4.
Idiomatic error handling. We describe how C++ can enhance MPI error handling via compile-time checking and exceptions in Section˜3.5.
These are the major aspects and principles to which a future C++ MPI interface should adhere. Rather than proposing a complete standardized interface, which risks repeating the shortcomings of the removed bindings, our primary goal is to define semantic guidelines and core conceptual interfaces that align idiomatic C++ with MPI. These concepts could serve as the foundation for a MPI standard side document. Standardizing concepts rather than concrete APIs offers the advantage of long-term maintainability and adaptability as both MPI and the C++ language evolve. A C++ interface could introduce additional quality-of-life features to enhance productivity and provide safer abstraction, but these often introduce overheads that conflict with MPI’s performance-portability while increasing the complexity of a specification. Still, we outline major ideas and provide suggestions on how they could be aligned to our previously introduced design considerations. Particularly, we focus on serialization of more complex data types (Section˜4.1), first steps towards automatic life cycle handling of more complex MPI data types with direct mappings to C++ (Section˜4.2) and how to increase MPI programmer’s productivity by a sane set of defaults (Section˜4.3).
In Section˜5, we conclude our discussion with suggestions on how the MPI standard could support language interface designers (not limited to C++) by carefully extending the MPI specification.
@@ 2 Existing C++ Interfaces over MPI
Since the removal of the “official” C++ MPI interface, a large number of third-party library interfaces have emerged, joined by efforts to bridge the gap between MPI implementations, C++, and high-level performance portability frameworks (e.g., Kokkos). We will briefly outline notable works and their distinctive features in the following, as they serve as inspirations and proofs of concept for the technical discussion in the rest of this paper.
@@
RWTH-MPI [9] is a C++ interface for MPI that supports contiguous STL containers as send/receive buffers and offers overloaded MPI procedures with automatic parameter inference. For custom types, it can auto-generate MPI data types using the PFR library [20]. Dynamic-size types are unsupported. While it covers MPI 4.0, its bindings largely mirror the C interface with limited added abstraction or safety.
@@
Kokkos Comm [2] utilizes modern C++ language features to provide more intuitive support of inter-node communication of Kokkos Views. Templates provide the needed flexibility to adapt to diverse data layouts and memory spaces.
@@ 3 High-Level Design Considerations
We start by defining how MPI objects map to C++ (Section˜3.1). Then we focus on how to model the actual data involved in communication, in terms of types (Section˜3.2) and the actual memory involved (Section˜3.3). We then discuss how combining C++’s ownership model and MPI to obtain additional memory safety (Section˜3.4) and how to handle errors in an idiomatic way (Section˜3.5).
@@ 3.1 Mapping the MPI Object Model to C++
For more than a decade, C++ MPI library designers have advocated for representing MPI objects as first-class C++ objects that act as proxies for the implementation-defined objects behind the handles [23].
It is worth noting that most MPI procedures operating on these objects can be expressed either as member functions or as free functions. Both choices are equally valid, and this decision is largely independent of the conceptual model discussed here. In this work, we present communication operations as member functions, without advocating for one approach over the other.
Modern C++ idioms—particularly RAII (Resource Acquisition Is Initialization) [25, 16.5][26, E.6] and move semantics—naturally support resource management: MPI objects can be constructed via MPI_*_create in class constructors and released via MPI_*_free in destructors, with ownership safely transferable through move operations. However, the original MPI design deviates from this model because of its reliance on global state and implicitly managed global objects—an approach discouraged by the C++ Core Guidelines [26, I.3,I.22]. For example, global communicators such as MPI_COMM_WORLD are automatically initialized via MPI_Init and cannot be explicitly created or freed. Further, MPI prohibits reinitialization after MPI_Finalize, limiting modular or library-based usage patterns.
@@ 3.1.1 Sessions, Groups, and Communicators
MPI 4.0 addressed some of these limitations by introducing the Session Model, an alternative to the traditional World Model for process management. Although originally designed to support better isolation in multi-threaded and multi-component environments, the session model aligns well with modern C++ object-oriented design. In this model, no global communicators are predefined; instead, communicators must be explicitly created from process sets via user code. This makes ownership explicit and fully under user control, enabling clean integration with RAII-based designs as shown in Fig.˜1. For example, communicators and groups are constructed and destroyed as part of their enclosing scope.
@@
The concept of constness is important for both the compiler and, more crucially, libraries, as it indicates variable mutability and restricts modifications in certain contexts. This applies straightforwardly to the data being communicated: For instance, a receiving buffer cannot be const, which already enhances the program’s safety, which we discuss in detail in Section˜3.3. In a less obvious way, constness can also be applied to library objects of a MPI C++ interface, such as communicators, data types, and requests and procedure calls on them, making it even more important to discuss here.
@@
We want to illustrate this for MPI communicators. Most procedure calls leave a communicator in the same state as before, which would naively allow them to be specified as const. Conversely, one could argue that posting a message is not the same as not having posted it; in the former case, the message can be received on the other side, while in the latter, a receive operation may hang. We think this is a strong argument to mark communication calls as non-const. But this can still be challenged for procedures that send and receive in a single call (such as send_recv or collectives).  [^1] Note that this is a fundamental problem, not exclusive to const-qualified member functions; if the procedures are free functions the questions will still stand regarding the constness of the communicator argument parameter.
To resolve these cases, we invoke a modern interpretation (post C++11): Since the communicator is likely undergoing internal mutation during any non-trivial operation, it is reasonable to conclude that most communication procedures should be marked as non-const at the communicator level.
@@
auto size() const; // most likely can be marked const
// in a reasonable implementation
@@
If we adhere to the conclusion that communication procedures should not be marked const, an important and somewhat unexpected result follows: a communicator class should not have a duplicate mechanism that is const on the original communicator. Incidentally, there will be no canonical copy constructor (taking original communicator as a constant reference). This is consistent, since even communicators duplicated from each other cannot receive or complete operations initiated in another communicator. Thus, two communicators are never copies of each other; they are at most alternative virtual fabrics for subsequent communication. Fig.˜2 summarizes our discussions about the communicator class which also align with the interpretation of the B.MPI3 [7] interface.
@@ 3.2 Modeling and Mapping Types
Using template-metaprogramming, a C++ MPI interface can therefore deduce an MPI type directly from a data buffer (as defined in Section˜3.3), in case its underlying value_type is fundamental. This approach is implemented by all major MPI C++ bindings [3, 7, 13, 27]. More complex types require explicit creation and a subsequent commit step, and have to be freed before MPI is finalized. To enable proper cleanup we again use RAII and represent data types as mpi::datatype objects, which support move construction and assignment and free the data type when the destructor is called.
To prevent users from using uncommitted data types in communication, we propose to encode the commit information as part of the type, similar to the approach of rsmpi [24]. The function mpi::commit takes an rvalue mpi::datatype object, and converts it to a mpi::committed_datatype as shown in Fig.˜3. Accordingly, communication calls only accept data type objects of this type.
@@
std::array<int, 3> b;
double c;
char d;
@@
These types can be safely supported by automatically constructing the corresponding MPI datatype using a compile-time reflection mechanism to call the correct MPI type constructors. This ensures that the C++ type is compatible with MPI’s type system and enables safe type handling, even on heterogeneous systems. Although C++ currently lacks native language support for such reflection, concrete proposals are in place to introduce compile-time type reflection in C++26 [22]. In the meantime, third-party libraries such as Boost.PFR [20] can be used to implement this feature, although they have some limitations, for example for types using inheritance or private members.
Beyond fundamental and trivially copyable types, there remains a broad class of other (potentially user-defined) C++ types that do not have a direct or automatically derivable mapping to an MPI data type. For these types, the correspondence between memory layout and type semantics cannot be safely inferred by the prospective C++ MPI interface. Consequently, it becomes the user’s responsibility to explicitly define how the type is laid out in memory and how that layout corresponds to an appropriate MPI datatype. While the handling of fundamental and trivially copyable types can be automated to reduce programming errors and ease MPI development, a prerequisite for such automation is a clear conceptual model of the data involved in communication. The core abstraction in this model is the data buffer, which we introduce in the following Section˜3.3.
@@ 3.3 Modeling Memory Involved in MPI Communication
Particularly when using standard library containers such as std::vector<T> with the MPI C API, users have to access underlying raw pointers and sizes of the containers and pass them to MPI individually, while the actual vector object already comprises the whole send data context. This is both non-ergonomic to use and conflicts with the C++ Core Guidelines [26, I.13]. Most existing MPI C++ libraries therefore introduce support for a subset of standard library containers [3, 7, 9, 13, 27] as their core abstraction feature. However, this has certain shortcomings. First, this is often defined ad-hoc only including a fixed set of containers [9, 13]. Second, design decisions within the approaches of supporting (standard library) containers introduce hidden, additional overheads through memory allocations or additional MPI calls [3, 13, 27]. For example, in an MPI_Recv call, Boost.MPI resizes the container into which it receives the data to match the size of the incoming message. If the user does not want this, they must instead pass raw pointers again. The KaMPIng library internally invokes MPI_Probe when the user issues a receive operation without a count argument. Such unexpected overheads are not desired from the perspective of a standardized C++ interface.
@@
For each send or receive buffer involved in a communication call, MPI’s C API expects a pointer to a contiguous memory location, a count argument and an MPI data type. C++ already provides type information, and automatically matching types is crucial for type safety. A high-level standardizable C++ interface, can use type introspection to provide such type safety, which we detail in Section˜3.2. For now, let us assume that we can always deduce an MPI data type from a provided container. We call this basic building block of MPI communication (memory location, count and type) a data buffer. In the following, we describe a minimal interface of data buffer objects in terms of C++ concepts.
@@
  • std::ranges::contiguous_range, that is, its underlying memory has to be contiguous in memory. This captures MPI’s requirement that data which is communicated must reside in sequential storage  [^2] An exception to this requirement is the usage of MPI_BOTTOM which relies on absolute memory addresses to describe data and should be treated separately..
@@
An example of how these constraints can be expressed in plain C++ is given in Fig.˜4, which defines the concept mpi::DataBuffer in terms of the three concepts described above. A convenient property of this data buffer definition is that many C++ standard library containers like std::vector or std::array storing fundamental C++ types such as char, int, double, ... already satisfy the mpi::DataBuffer concept. This also holds for lightweight, non-owning views such as std::span. Hence, they can be directly used in an MPI C++ interface adhering to these principles.
For more complex cases, for example, to communicate a container of a custom C++ datatype which is trivially_copyable, we first have to construct and commit such a datatype (see Section˜3.2) and then pass it to the data buffer. For communicating data using complex derived datatypes without a one-to-one mapping to the underlying C++ type, we not only have to construct and commit an MPI datatype but also pass the number of corresponding elements to the data buffer. For these more complex cases, there is no direct C++ standard library support but an MPI C++ interface could easily provide a lightweight adapter with multiple constructors not only accepting a contiguous std::range, but also a count and a mpi::committed_datatype parameter.
This also makes this interface extensible to support more complex containers from third-party libraries like Kokkos::View or thrust::device_vector, backed by memory located on accelerators such as GPUs. Therefore the data buffer concept naturally extends to the notion of memory allocation kinds [18], providing direct support for accelerator-aware MPI on C++ containers backed memory located on GPUs or other accelerators.
Figure˜5 illustrates exemplary communication calls showcasing the different cases discussed above. Note that we do not suggest a concrete API or function signatures but rather want to show the underlying concept of using data buffers as an abstract description for communication data.
Orthogonal to the mpi::DataBuffer concept presented so far, we additionally have to model that data buffers describing send data have to be readable and data buffers for receive data have to be writable. This can be expressed using the (built-in) concepts, std::ranges::in/output_range (see Fig.˜4).
@@
std::vector<MyType> v = ... // static custom type defined in Fig.˜3
@@
Figure 5: Exemplary communication calls using the data buffer concept when MPI procedures are implemented as member functions.
Extending the Data Buffer Concept.
In collective communication calls with varying counts (such as MPI_Alltoallv), MPI requires not just a single count parameter but a separate count for each MPI process (aka rank). Additionally, displacements can be specified for each rank. To reflect these semantic changes, we have to extend the mpi::DataBuffer concept as previously defined (Fig.˜4) and define irregular mpi::DataBuffers. These buffers are still based on the concept std::ranges::contiguous and have to expose MPI datatype information. Additionally, they also have to expose size_v() and displacements() functions, which return the respective counts and displacement data. For brevity, we refrain from providing an example definition of this additional concept. Again, simple lightweight adapters can be provided for standard library and similar custom containers. Similarly, scalar arguments (such as a single int or bool in a reduction) can be easily supported via lightweight adapters, which implicitly return size $1$ and implement the data buffer concept (Fig.˜5).